OpenStudy (anonymous):

Choose the equation of the line passing through the point (6, -2) and parallel to y = x + 7.

OpenStudy (anonymous):

slope will remain 1 y - y1 = m (x - x1)

OpenStudy (lalaly):

parallel means they have the same slope so slope=1 y=mx+c substitute the point (6,-2) im x and y -2=(1)(6)+c -2=6+c -8=c so the equation of the line is y=x-8

OpenStudy (anonymous):

oh thank you so much. i understand how to do it now.